A secretary makes 2 errors per page on the average. What is the probability that on the next 

page she makes; 4 or more errors and no error?


1
Expert's answer
2022-06-01T13:10:43-0400

Average error per page "\lambda=2"

Let X denote the number of Errors: "X\sim Po(\lambda)"


(a) 

"P(X\ge4)=1-P(X=0)-P(X=1)"




"-P(X=2)-P(X=3)"

"=1-\dfrac{e^{-2}(2)^0}{0!}-\dfrac{e^{-2}(2)^1}{1!}-\dfrac{e^{-2}(2)^2}{2!}"

"-\dfrac{e^{-2}(2)^3}{3!}=0.14287"


(b) 

"P(X=0)=\dfrac{e^{-2}(2)^0}{0!}=0.13534"